The Advanced Photon Source
a U.S. Department of Energy Office of Science User Facility

APS-UES Meeting Minutes - 12/05/2016

Attendees: Bruce Glagola, Bruno Fieramosca, Clay White, Nena Moonier, Wendy VanWingeren, Shane Flood, Dean Wyncott, and Patti Pedergnana

Guest: Julie Cross

Bruce Glagola

  • Performance appraisals were released on Thursday of last week.  The FCs need to review their appraisals and submit their electronic signatures.
  • Bruno will attend the Shutdown Planning meeting on Wednesday in Bruce’s absence.  Bruce will be out of town Wednesday through Saturday.
  • There are memos out about effort recording and holiday payrolls.  December 22nd is the last day of work before holidays, then the lab is closed December 27th-29th, followed by the New Year holiday.  Bruce has to approve the week ending 12/18 and 12/25 by 12/19. 
  • Clay and Jonathon Baldwin have picked up laboratory L1119 DEEI inspections that needs to be done by the end of January.  He asked Bruno and Dean to assist.
  • The UES group lunch will be scheduled during shutdown.

Dean Wyncott

  • The supply side DI water at 8-ID sprung a heavy leak on a breaker panel.  This wound up giving about 30 minutes of downtime at the beamline.
  • 35-ID is not going to put their mirror box back together until shutdown.  The HP type C survey will be done on the first day of January 2017’s start up.
  • 33-ID will put a cryo-pump on top of the A station.  Central Shops should be building a deck for the beamline.
  • Sector 22’s drain line did not line up to the drain.  Water was pouring all over the beamline.  Dean contacted FMS and found tarps to help cover the beamline and control area.  FMS rerouted the drain pipes to the drain.
  • Chris Roehrig said there was a beam dump on Friday and his 2-ID front end shutter did not close.
  • Dean asked Julie Cross about an issue with putting equipment on station roofs.  Dean has worked with John Sidarous regarding the weight loads on roofs and Sidarous said that the roof supports can withstand the loads put on the roofs, but not the areas in between the roof supports.

Patti Pedergnana

  • Changyong Park at 16-BM-D wants to replace the station doors and possibly the pneumatics too.  Andy Stevens, Tim Clute, Joe Budz, and Anne Boron will meet with HP-CAT personnel on Monday, December, 12th to develop a scope of work to send to Tecknit.
  • MOM removed the ODH monitor at 13-ID-A for shipment offsite.
  • The roofers were out last week to repair the 434 A hallway ceiling leak.  They found a number of places leaking in the roof and should have repaired them.
  • The XSD-TRR Three-Year Safety Review will be rescheduled due to the APS All Hands Meeting being held on the same day at the same time.

Shane Flood

  • Sector 22 wants to tape the crack on the 22-ID-A station roof with lead tape.  Julie Cross said that they should use a different type of tape as the leaded tape adds a hazard that people will be walking on.

Clay White

  • 10-ID wants to change the limits on their undulator which is currently at 13 mm.  John Grimmer was consulted on the matter and will contact Clay soon.

Bruno Fieramosca

  • 5-BM and 5-ID still have to be consulted on their water needs during shutdown.  Bruce asked Dean to double check with his area again, he had a few groups that had not responded.
  • Bruno requested that we get the 431 Floor Coordinator office back.

Julie Cross

  • 2-ID’s shutter removal has been approved.  There are hard stops on the mirror that need to be tagged.  Dean will tag the hard stops today or next Monday’s machine studies.  Julie and Dean will look at the stops in the 2-ID-A station next Monday.
  • Julie asked if the FC group needs help collecting LN2 system drawings from the beamlines.  There will be training on the LNDS given to a few of the MOM group technicians.